# VGLD: Visually-Guided Linguistic Disambiguation for Monocular Depth Scale Recovery

### v2

@@ 3.3 Loss Function
4 Experiments
4.1 Experimental Settings
Dataset. We primarily train on two datasets: NYUv2[43] and KITTI[14], representing indoor and outdoor scenes, respectively. NYUv2 contains images with a resolution of 480x640, with depth values ranging from 0 to 10 meters. In accordance with the official dataset split[25], we use 24,231 image-depth pairs for training and 654 image-depth pairs for testing. KITTI is an outdoor dataset collected from equipment mounted on a moving vehicle, with depth values ranging from 0 to 80 meters. Following KBCrop[46], all RGB images and depth maps are cropped to a resolution of 1216x352. We adopt the Eigen split[9], which includes 23,158 training images and 652 test images, to train and evaluate our method. Additionally, we report zero-shot generalization results on SUN-RGBD[44], which includes 5,050 test images, DIML Indoor[8], which contains 503 validation images and DDAD[15], which contains 3950 validation images.
Relative Depth Models. We use MiDaS 3.1[6] with the dpt_swin2_large_384 model (213M parameters), MiDaS 3.0[39] with the dpt_large_384 model (123M parameters), DepthAnything[53] with DepthAnything-Small model (24.8M parameters), and DepthAnything v2[54] with DepthAnything-V2-Small model (24.8M parameters). Additionally, we apply the Least Squares and Levenberg-Marquardt algorithms, which predict a global scale and offset based on the corresponding ground-truth depth maps, scaling relative depth accordingly, as detailed in Tables 1, 2, 7, 8.
The Proposed Models. For clarity, we denote the proposed models as VGLD-{dataset}-{method}. The dataset refers to the training datasets, which include "N" for NYUv2, "K" for KITTI, and "NK" for both NYUv2 and KITTI. The method refers to the type of embeddings used: "T" for text embeddings only, "I" for visual embeddings only, and "TCI" for both text and visual embeddings (i.e., Fusion Embeddings, as shown in Figure 2).
Training details. The proposed VGLD is implemented in PyTorch. We use the Adam optimizer with parameters $(\beta_{1},\beta_{2},\text{wd})=(0.9,0.999,0.001)$ and a learning rate of $3\times 10^{-4}$. All models are trained for 24 epochs on a single NVIDIA RTX 3090 GPU with 24GB of memory. The batch size is set to 6, and the total training time for each model is approximately 19 to 22 hours.
[…]
ZoeDepth[5] robust depth estimation $\dagger$ 0.077 0.277 0.033 0.953 0.995 0.999
ZeroDepth[16] 0.074 0.269 0.103 0.954 0.995 1.000
Metric3Dv2[17] 0.047 0.183 0.020 0.989 0.998 1.000
MiDas-1[6] Least Squares 0.121 0.073 0.388 0.338 0.068 0.866 0.959 0.978
@@ 4.1 Experimental Settings
MiDas-2[39] Least Squares 0.130 0.085 0.421 0.286 0.066 0.845 0.956 0.980
@@
DAV2-vits[54] Least Squares 0.122 0.074 0.392 0.362 0.070 0.866 0.959 0.977
@@
DAV1-vits[53] Least Squares 0.121 0.075 0.397 0.327 0.067 0.863 0.959 0.979
@@
Table 1: Quantitative Depth Comparison on the NYUV2 Dataset. The yellow regions represent the results obtained using least-squares-type mathematical fitting methods. $\dagger$ denotes the results of certain state-of-the-art (SOTA) absolute scale estimation models. $\ddagger$ In the RDE Model column, MiDas-1 refers to the model using Midas-V3.1-dpt_swin2_large_384, MiDas-2 refers to the model using Midas-V3.0-dpt_large_384, DAV2-vits refers to the model using Depth-Anything-V2-Small, and DAV1-vits refers to the model using Depth-Anything-V1-Small. Best results are in bold, second best are underlined.
RDE Model Method Lower is better Higher is better
ZoeDepth[5] robust depth estimation 0.054 2.281 0.082 0.971 0.996 0.999
ZeroDepth[16] 0.053 2.087 0.083 0.968 0.995 0.999
Metric3Dv2[17] 0.044 1.985 0.064 0.985 0.998 0.999
MiDas-1[6] Least Squares 0.333 2.094 6.901 1.731 0.293 0.408 0.790 0.879
@@
MiDas-2[39] Least Squares 0.336 2.172 6.925 1.658 0.283 0.421 0.778 0.876
@@
DAV2-vits[54] Least Squares 0.330 2.053 6.737 1.729 0.292 0.423 0.790 0.877
@@
DAV1-vits[53] Least Squares 0.331 2.078 6.772 1.714 0.291 0.423 0.786 0.875
@@
Table 2: Quantitative Depth Comparison on the KITTI Dataset. The yellow regions represent the results obtained using least-squares-type mathematical fitting methods. Best results are in bold, second best are underlined.
Figure 5: Sensitivity to variations in linguistic descriptions on the NYUv2 dataset. Depth maps show that the RSA method is susceptible to different textual inputs, resulting in inconsistent estimations, especially in the regions highlighted by orange boxes. In contrast, our VGLD method yields more stable predictions. Warmer colors (red) indicate closer distances, while cooler colors (blue) indicate farther distances.
4.2 Experimental Results
Quantitative results. We present the results on NYUv2 in Table 1 and on KITTI in Table 2. Our approach surpasses RSA[59] in all evaluation metrics and achieves performance close to that of scaling with ground-truth values, as highlighted in yellow in the all quantitative tables. Following [39], we perform linear regression to find the scale and shift that minimize the least-squares error between the ground-truth and predicted metric depth. Additionally, we employ the Levenberg-Marquardt algorithm for more accurate nonlinear fitting. We trained separate VGLD models for each dataset and also a unified VGLD model combining NYUv2 and KITTI, denoted as VGLD-NK-{method}. The VGLD model trained on a single dataset shows slightly better performance compared to the unified model, benefiting from the precise routing functionality of the DRM module, the performance gap between the single-dataset and unified models is minimal. Furthermore, VGLD models using visual embeddings (VGLD-{dataset}-I) consistently outperform those using only text embeddings (VGLD-{dataset}-T), validating the effectiveness of using visual rather than purely linguistic cues for scale prediction. Building on this, we combine both visual and text embeddings (VGLD-{dataset}-TCI), enabling visual embeddings to guide text semantic alignment and addressing the visually-guided linguistic disambiguation problem.
Qualitative comparisons. We present comparison examples of VGLD and baseline methods on the NYUv2 and KITTI datasets in Figures 1 and 2, respectively. The error maps display the absolute relative error, where the overall brightness of the error maps clearly indicates the performance of our method. Notably, our approach achieves performance very close to that of the Levenberg-Marquardt fitting (LM Fit) across different scenes, demonstrating robust metric depth scale recovery. In contrast to the fixed scale and shift estimates produced by RSA, VGLD significantly improves the accuracy of depth predictions, with darker error maps indicating reduced error. Note: All qualitative comparison results in the VGLD section are inferred from the VGLD-NK-TCI method, where the RDE model used is DAV1-vits.
Sensitivity to Variations in Linguistic Descriptions. A single image can be described using multiple textual expressions. To investigate how linguistic variation affects metric depth scale recovery, we evaluate the influence of different textual inputs on VGLD’s performance. We observe that while the RSA method—relying solely on textual descriptions—is highly sensitive to phrasing, VGLD exhibits significantly greater robustness, consistently producing stable predictions for scale and shift. Moreover, VGLD often outperforms RSA across evaluation metrics, highlighting its ability to provide reliable scalars estimations. Figures 5 and 6 present qualitative comparisons on NYU and KITTI under three distinct text prompts. The corresponding quantitative results are provided in Tables 3 and 4, while Tables 5 and 6 list the three textual descriptions used for each image, as detailed in the Supplementary Material (Section 7)..
Figure 6: Sensitivity to variations in linguistic descriptions on the KITTI dataset. Depth maps show that the RSA method is susceptible to different textual inputs, resulting in inconsistent estimations, especially in the regions highlighted by orange boxes. In contrast, our VGLD method yields more stable predictions. Warmer colors (red) indicate closer distances, while cooler colors (blue) indicate farther distances.
5 Conclusion
We propose an enhanced framework, VGLD, that leverages linguistic descriptions as input and utilizes high-level semantic information from images to address the issue of linguistic ambiguity. VGLD effectively harnesses pre-trained CLIP Text Encoder and Image Encoder to extract rich semantic information from both text and images, and jointly employs these features through an MLP-based model to map them into shift and scale parameters, robustly converting relative depth maps into metric-depth predictions. Extensive experiments on indoor and outdoor benchmark datasets, along with four pre-trained relative depth models, demonstrate VGLD’s strong robustness, validating that vision can correctly guide language and resolve textual ambiguities during inference, enabling accurate depth scale recovery without ground truth values. We also present a unified model, VGLD-NK-{method}, which is trained across both indoor and outdoor datasets. Thanks to the design of the Domain Router Mechanism (DRM), the VGLD model can easily generalize across different scene compositions, highlighting the importance of the DRM module. Finally, a zero-shot generalization experiment shows that, even on unseen data domains, VGLD can effectively guide linguistic inputs without additional training. While the linguistic modality is invariant to common visual variations, the sole reliance on language modalities can still struggle to fully and accurately describe a specific image. In comparison to additional modalities (e.g., IMU, LiDAR sensors), using the image modality to complement the required semantic information for the task is a convenient and effective approach.
5.1 Limitations and future work.
Although linguistic-based scale recovery under visually-guided methods is highly robust, VGLD is still influenced by language modality. For different descriptions of the same image, the VGLD model may output inconsistent results (albeit with small error margins), especially when incorrect descriptions are used (e.g., describing an indoor scene as "a photo of a narrow street."). To address this issue, one feasible approach could be to further match the similarity between the language and image modalities, effectively excluding erroneous image descriptions. Future work could expand the image modality-assisted features of VGLD to enable more robust and fine-grained scale estimation, as well as enhance the model’s ability to handle malicious attacks in text descriptions.
Supplementary Material
6 Evaluation Metrics
We evaluate our approach using the standard five error metrics and three accuracy metrics commonly adopted in prior works[42]. Specifically, the error metrics include absolute mean relative error (Abs Rel), square relative error (sq_rel), log error($\log_{10}$), root mean squared error (RMSE), and its logarithmic variant ($\text{RMSE}_{\text{log}}$). The accuracy metrics are based on the percentage of inlier pixels ($\delta$) within three thresholds: $\delta_{1}<1.25$, $\delta_{2}<1.25^{2}$, and $\delta_{3}<1.25^{3}$.
7 Quantitative Results on Sensitivity to Linguistic Description Variations
As shown in Tables 3 and 4, We quantitatively evaluated the inference results and sensitivity of the VGLD model to variations in linguistic descriptions. For both indoor and outdoor datasets, three images were used, with each image paired with three distinct textual descriptions. The corresponding visualization figures are provided in Figure 5 and 6 . And the specific textual descriptions are provided in Tables 5 and 6.
@@ 7 Quantitative Results on Sensitivity to Linguistic Description Variations
Table 3: Quantitative results on the NYUv2 dataset comparing VGLD and RSA in response to different textual descriptions. The yellow regions represent scalars values fitted using the Levenberg-Marquardt method. Best results are in bold.
1 Text-1 RSA[59] 0.097 4.060 0.926 1.005 1.011
VGLD 0.075 3.562 0.949 1.004 1.001
Text-2 RSA[59] 0.084 4.251 0.923 1.007 1.010
VGLD 0.077 3.067 0.949 1.004 1.010
Text-3 RSA[59] 0.072 3.140 0.952 1.004 1.010
VGLD 0.068 3.088 0.951 1.004 1.003 1.010 1.010
2 Text-1 RSA[59] 0.108 2.327 0.905 1.009 1.014
VGLD 0.063 1.887 0.984 1.135 1.015
Text-2 RSA[59] 0.281 5.341 0.537 1.006 1.013
VGLD 0.099 2.144 0.915 1.010 1.014
Text-3 RSA[59] 0.109 2.157 0.906 1.011 1.014
VGLD 0.073 1.861 0.952 1.011 1.009 1.015 1.017
3 Text-1 RSA[59] 0.268 6.526 0.740 1.004 1.009
VGLD 0.119 2.516 0.919 1.009 1.010
Text-2 RSA[59] 0.171 4.479 0.854 1.005 1.010
VGLD 0.077 2.287 0.942 1.008 1.010
Text-3 RSA[59] 0.081 2.421 0.938 1.007 1.010
VGLD 0.062 2.236 0.953 1.009 1.008 1.011 1.011
Table 4: Quantitative results on the KITTI dataset comparing VGLD and RSA in response to different textual descriptions. The yellow regions represent scalars values fitted using the Levenberg-Marquardt method. Best results are in bold.
@@
8 Zero-shot Generalization
Benefiting from the smaller domain gap of language descriptions across diverse scenes[58, 59] and the ability of corresponding images to accurately indicate domain context, we conduct a zero-shot transfer experiment to demonstrate the generalization capability of VGLD. We evaluate the models on the SUN-RGBD[44] , DIML Indoor[8], and DDAD[15] datasets without any fine-tuning. As shown in Figures 7, 8 ,9 (qualitative results) and Tables 7, 8, 9 (quantitative results), VGLD consistently outperforms baseline methods and produces results that closely match those fitted by the LM method. This demonstrates that, under visual guidance, VGLD maintains stable scalars estimation and exhibits enhanced generalization capabilities. Note that all zero-shot experiments are conducted using the VGLD-NK-TCI model built upon the DAV1-vits RDE backbone.

### v3

@@ Introduction
Figure 4: Sensitivity to variations in linguistic descriptions on the KITTI dataset. Similar to Figure 3, we focus on the differences within the orange boxes across the three textual inputs. Note that we use LM fitting results instead of the ground-truth depth map for visualization, as the KITTI ground-truth data is too sparse to yield meaningful visual comparisons. Warmer colors (red) indicate closer distances, while cooler colors (blue) indicate farther distances.
Experiments
Experimental Settings
Dataset. We primarily train on two datasets: NYUv2(Silberman et al. 2012) and KITTI(Geiger, Lenz, and Urtasun 2012), representing indoor and outdoor scenes, respectively. NYUv2 contains images with a resolution of 480x640, with depth values ranging from 0 to 10 meters. In accordance with the official dataset split(Lee et al. 2019), we use 24,231 image-depth pairs for training and 654 image-depth pairs for testing. KITTI is an outdoor dataset collected from equipment mounted on a moving vehicle, with depth values ranging from 0 to 80 meters. Following KBCrop(Uhrig et al. 2017), all RGB images and depth maps are cropped to a resolution of 1216x352. We adopt the Eigen split(Eigen, Puhrsch, and Fergus 2014), which includes 23,158 training images and 652 test images, to train and evaluate our method. Additionally, we report zero-shot generalization results on SUNRGBD(Song et al. 2015), which includes 5,050 test images, DIML Indoor(Cho et al. 2021), which contains 503 validation images and DDAD(Guizilini et al. 2020), which contains 3950 validation images.
Relative Depth Models. We use MiDaS 3.1(Reiner et al. 2023) with the dpt_swin2_large_384 model (213M parameters), MiDaS 3.0(Ranftl et al. 2020) with the dpt_large_384 model (123M parameters), DepthAnything(Yang et al. 2024a) with DepthAnything-Small model (24.8M parameters), and DepthAnything v2(Yang et al. 2024b) with DepthAnything-V2-Small model (24.8M parameters).
The Proposed Models. For clarity, we denote the proposed models as VGLD-{dataset}-{method}. The {dataset} refers to the training datasets, which include ”N” for NYUv2, ”K” for KITTI, and ”NK” for both NYUv2 and KITTI. The {method} refers to the type of embeddings used: ”T” for text embeddings only, ”I” for visual embeddings only, and ”TCI” for both text and visual embeddings (i.e., Fusion Embeddings, as shown in Figure 2).
Experimental Results
Quantitative results. We present the results on the NYUv2 and KITTI datasets in Table 1. (More detailed quantitative results are provided in Table 5 and Table 6 in the Supplementary Material.). Our approach consistently outperforms RSA(Zeng et al. 2024b) across all evaluation metrics and achieves performance comparable to scale recovery using ground-truth depths, as indicated in the Least Squares and Levenberg-Marquardt sections of the quantitative tables. The quantitative results show that models trained on a single dataset (VGLD-N or VGLD-K) perform slightly better within their respective domains compared to the unified model VGLD-NK. For example, VGLD-N/K-TCI with DAV2-ViTS as the RDE model achieves the best performance across all three evaluation metrics reported in the Table 1. Thanks to the precise routing capability of the DRM module, the performance gap between the single-dataset and unified models remains marginal, highlighting the strong cross-domain generalization ability of the unified VGLD-NK model. For example, based on DAV2-ViTS, the VGLD-N-TCI model achieves an AbsRel of 0.125 on NYUv2, and the VGLD-K-TCI model achieves 0.152 on KITTI. The unified VGLD-NK-TCI model obtains AbsRel scores of 0.127 and 0.153 on NYUv2 and KITTI, respectively, representing decreases of less than 1.58% and 0.65%.
Furthermore, models utilizing visual embeddings (VGLD-XX-I) consistently outperform those relying solely on textual embeddings (VGLD-XX-T), validating the effectiveness of visual cues for scale prediction over purely linguistic prompts. For example, based on DAV1-ViTS, the VGLD-NK-T model achieves AbsRel scores of 0.142 and 0.148 on NYUv2 and KITTI, respectively. In comparison, VGLD-NK-I achieves AbsRel scores of 0.114 and 0.142 on the same datasets, corresponding to improvements of 24.5% and 4.2%, respectively. Building on this, we combine both visual and textual embeddings (VGLD-XX-TCI), allowing visual features to guide the semantic alignment of textual inputs. This integration yields modest but meaningful improvements, thereby effectively addressing the challenge of visually grounded linguistic disambiguation.
Notably, the improvement of VGLD-XX-TCI over VGLD-XX-T is less pronounced on KITTI compared to NYUv2. We attribute this to the lower variance in outdoor scene descriptions in KITTI, whereas indoor scenes in NYUv2 exhibit much greater diversity—such as bathrooms, kitchens, classrooms… This higher variability in textual descriptions benefits the model by providing richer cues for more accurate estimation of scene-specific scaling parameters.
For completeness, the Supplementary Material presents more extensive quantitative results and qualitative comparisons, including those from the zero-shot evaluation setting.
Sensitivity to Variations in Linguistic Descriptions. A single image can be described using multiple textual expressions. To investigate how linguistic variation affects metric depth scale recovery, we evaluate the influence of different textual inputs on VGLD’s performance. Figures 3 and 4 present qualitative comparisons on NYU and KITTI under three distinct text prompts. We observe that while the RSA method—relying solely on textual descriptions—is highly sensitive to phrasing, VGLD demonstrates significantly greater robustness, consistently producing stable predictions for both scale and shift. This is most evident in the third image of Figure 3: RSA accurately recovers the depth when paired with Text-3 (whose prediction closely matches the ground truth), but exhibits substantial errors with Text-1 and Text-2. In contrast, VGLD achieves stable and accurate scale recovery across all three descriptions (Text-1 to Text-3). Moreover, VGLD often outperforms RSA across evaluation metrics, further highlighting its ability to provide reliable scalar estimations. The corresponding quantitative results are provided in the Supplementary Material, along with the three textual descriptions used for each image.
Ablation Study
Effect of the DRM. As shown in Table 2, we conduct ablation studies on the Domain Router Mechanism (DRM). The results demonstrate that incorporating the DRM consistently improves the overall performance of VGLD across all four backbone models and significantly enhances its cross-domain generalization capability. The ablation studies are conducted based on the VGLD-XX-TCI model.
Models Method NYU KITTI
[…]
MiDas-1 w/o DRM 0.128 0.415 0.757 0.136 3.636 0.855
with DRM 0.120 0.414 0.863 0.121 3.559 0.874
MiDas-2 w/o DRM 0.158 0.513 0.740 0.198 4.987 0.728
with DRM 0.151 0.513 0.780 0.185 4.804 0.737
DAV2-vits w/o DRM 0.135 0.459 0.798 0.163 4.060 0.767
with DRM 0.127 0.434 0.835 0.153 3.980 0.772
DAV1-vits w/o DRM 0.122 0.437 0.847 0.145 4.327 0.748
with DRM 0.112 0.392 0.883 0.136 4.008 0.816
Table 2: Performance comparison on NYU and KITTI datasets with and w/o(without) DRM. Best results are in bold.
Effect of the LM loss. To investigate the effect of different weights of LM loss $\mathcal{L}_{lm}$ on model training, we vary the value of $\beta$ in equation 5 and train the VGLD-NK-TCI model based on the DAV1-vits RDE backbone. Evaluation on both the NYUv2 and KITTI datasets shown in Table 3 that the model achieves the best performance when $\beta=0.1$. Compared to completely removing the $\beta$ term ($\beta=0$), the model achieves a 2.7% improvement in AbsRel on NYUv2 and a significantly larger gain of approximately 20.5% on KITTI. This demonstrates the effectiveness of the $\mathcal{L}_{\text{lm}}$ constraint, particularly in more open outdoor environments, where stronger guidance is needed to stabilize the training trajectory.
$\beta$ NYU KITTI
[…]
0 0.115 0.403 0.874 0.164 4.856 0.781
0.001 0.116 0.413 0.869 0.161 4.204 0.779
0.01 0.113 0.399 0.879 0.146 4.010 0.791
0.1 0.112 0.392 0.883 0.136 4.008 0.816
1 0.115 0.397 0.868 0.162 4.701 0.778
Computational Complexity. As shown in Table 4, we present a comparison of model parameters and inference times between VGLD and RSA to quantify the computational resources required. All evaluations were conducted on a single NVIDIA RTX 3090 (24GB). This experiment is conducted using the DAV1-vits RDE backbone. The results indicate that the scalar predictor in VGLD is more lightweight and efficient compared to that of RSA. However, VGLD additionally incorporates a CLIP image encoder, which introduces an extra 14ms of inference time compared to RSA. Despite this overhead, VGLD offers a favorable trade-off: it achieves a 32.1%(Ref. to Tabel 1) improvement in Abs Rel on NYUv2 with an inference time of just 14.08ms increases and a modest parameters, making it a practical and efficient choice.
Components RSA VGLD (ours)
Params# Inf. Times Params# Inf. Times
DAV1-vits 24.78M 9.62ms 24.78M 9.62ms
CLIP Text Encoder 63.43M 13.61ms 63.43M 13.61ms
CLIP Image Encoder - - 86.19M 14.90ms
Scalars Predictor 1.49M 1.76ms 1.18M 0.94ms
Total 89.7M 24.99ms 175.58M 39.07ms
Increase / M (ms) - - 85.88M $\uparrow$ 14.08ms $\uparrow$
Table 4: Computational Complexity Analysis. As shown in the table, the increase in model parameters(Params#) and inference times(Inf. Times) of VGLD compared to the RSA model primarily stems from the additional CLIP Image Encoder component.
Conclusion
We presented VGLD, a novel framework for monocular depth scale recovery that performs Visually-Guided Linguistic Disambiguation. VGLD leverages high-level visual semantics to resolve inconsistencies in textual inputs, enabling stable and accurate scale prediction across diverse linguistic descriptions. By jointly encoding image and text via CLIP and predicting global transformation parameters with an MLP, VGLD transforms relative depth maps into metric estimates in a robust and consistent manner. Extensive evaluations on both indoor and outdoor benchmarks show that VGLD significantly reduces estimation variance under different captions and generalizes well across domains. Empowered by a Domain Router Mechanism, VGLD further supports universal deployment across scene types. Compared to sensor-based methods, VGLD offers a lightweight and effective alternative for reliable scale alignment.
Limitations and future work.
Although linguistic-based scale recovery under visually-guided methods is highly robust, VGLD is still influenced by language modality. For different descriptions of the same image, the VGLD model may output inconsistent results (albeit with small error margins), especially when incorrect descriptions are used (e.g., describing an indoor scene as ”a photo of a narrow street.”). To address this issue, one feasible approach could be to further match the similarity between the language and image modalities, effectively excluding erroneous image descriptions. Future work could expand the image modality-assisted features of VGLD to enable more robust and fine-grained scale estimation, as well as enhance the model’s ability to handle malicious attacks in text descriptions.
Appendix A Supplementary Material
Evaluation Metrics
We evaluate our approach using the standard five error metrics and three accuracy metrics commonly adopted in prior works(Shao et al. 2023). Specifically, the error metrics include absolute mean relative error (Abs Rel), square relative error (sq_rel), log error($\log_{10}$), root mean squared error (RMSE), and its logarithmic variant ($\text{RMSE}_{\text{log}}$). The accuracy metrics are based on the percentage of inlier pixels ($\delta$) within three thresholds: $\delta_{1}<1.25$, $\delta_{2}<1.25^{2}$, and $\delta_{3}<1.25^{3}$.
Training details
The proposed VGLD is implemented in PyTorch2.0.1+CUDA11.8. We use the Adam optimizer with parameters $(\beta_{1},\beta_{2},\text{wd})=(0.9,0.999,0.001)$ and a learning rate of $3\times 10^{-4}$. All models are trained for 24 epochs on a single NVIDIA RTX 3090 GPU with 24GB of memory, , running in Ubuntu 22.04. The batch size is set to 6, and the total training time for each model is approximately 19 to 22 hours.
Qualitative comparisons
We present comparison examples of VGLD and baseline methods on the NYUv2 and KITTI datasets in Figure 5 and Figure 6, respectively. The error maps display the absolute relative error, where the overall brightness of the error maps clearly indicates the performance of our method. Notably, our approach achieves performance very close to that of the Levenberg-Marquardt fitting (LM Fit) across different scenes, demonstrating robust metric depth scale recovery. In contrast to the fixed scale and shift estimates produced by RSA, VGLD significantly improves the accuracy of depth predictions, with darker error maps indicating reduced error. Note: All qualitative comparison results in the VGLD section are inferred from the VGLD-NK-TCI method, where the RDE model used is DAV1-vits.
Quantitative Results on Sensitivity to Linguistic Description Variations
As shown in Table 7 and Table 9, We quantitatively evaluated the inference results and sensitivity of the VGLD model to variations in linguistic descriptions. For both indoor and outdoor datasets, three images were used, with each image paired with three distinct textual descriptions. The corresponding visualization figures are provided in Figure 3 and Figure 4(within the main text).. And the specific textual descriptions are provided in Table 8 and Table 10.
Zero-shot Generalization
Benefiting from the smaller domain gap of language descriptions across diverse scenes(Zeng et al. 2024a, b) and the ability of corresponding images to accurately indicate domain context, we conduct a zero-shot transfer experiment to demonstrate the generalization capability of VGLD. We evaluate the models on the SUN-RGBD(Song et al. 2015) , DIML Indoor(Cho et al. 2021), and DDAD(Guizilini et al. 2020) datasets without any fine-tuning. As shown in Figure 7, Figure 8, Figure 9 (qualitative results) and Table 11, Table 12, Table 13 (quantitative results), VGLD consistently outperforms baseline methods and produces results that closely match those fitted by the LM method. This demonstrates that, under visual guidance, VGLD maintains stable scalars estimation and exhibits enhanced generalization capabilities. Note that all zero-shot experiments are conducted using the VGLD-NK-TCI model built upon the DAV1-vits RDE backbone.
Effect of the initial seeds
Prompts for Natural Text Generation
To generate natural and semantically rich image descriptions—rather than relying on fixed prompt templates—we employ two vision-language models: LLaVA-v1.6-Vicuna-7B and LLaVA-v1.6-Mistral-7B(Jia et al. 2022). To ensure diversity in the generated captions, each model is prompted using six distinct instruction templates. These prompt templates are listed in Table 14.
@@ Appendix A Supplementary Material
Models Methods Abs Rel $\downarrow$ sq_rel $\downarrow$ RMSE $\downarrow$ $\text{RMSE}_{\text{log}}$downarrow log10 $\downarrow$ D1 $\uparrow$ D2 $\uparrow$ D3 $\uparrow$
ZoeDept(Bhat et al. 2023)h robust depth estimation 0.077 0.277 0.033 0.953 0.995 0.999
ZeroDepth(Guizilini et al. 2023) 0.074 0.269 0.103 0.954 0.995 1.000
Metric3Dv2(Hu et al. 2024a) 0.047 0.183 0.020 0.989 0.998 1.000
MiDas-1(Reiner et al. 2023) Least Squares 0.121 0.073 0.388 0.338 0.068 0.866 0.959 0.978
@@
MiDas-2(Ranftl et al. 2020) Least Squares 0.130 0.085 0.421 0.286 0.066 0.845 0.956 0.980
@@
DAV2-vits(Yang et al. 2024b) Least Squares 0.122 0.074 0.392 0.362 0.070 0.866 0.959 0.977
@@
DAV1-vits(Yang et al. 2024a) Least Squares 0.121 0.075 0.397 0.327 0.067 0.863 0.959 0.979
@@
Table 5: More detailed quantitative depth comparison on the NYUv2 dataset. Best results are in bold, second best are underlined.
Models Methods Abs Rel $\downarrow$ sq_rel $\downarrow$ RMSE $\downarrow$ $\text{RMSE}_{\text{log}}$downarrow log10 $\downarrow$ D1 $\uparrow$ D2 $\uparrow$ D3 $\uparrow$
ZoeDepth(Bhat et al. 2023) robust depth estimation 0.054 2.281 0.082 0.971 0.996 0.999
ZeroDepth(Guizilini et al. 2023) 0.053 2.087 0.083 0.968 0.995 0.999
Metric3Dv2(Hu et al. 2024a) 0.044 1.985 0.064 0.985 0.998 0.999
MiDas-1(Reiner et al. 2023) Least Squares 0.333 2.094 6.901 1.731 0.293 0.408 0.790 0.879
@@
MiDas-2(Ranftl et al. 2020) Least Squares 0.336 2.172 6.925 1.658 0.283 0.421 0.778 0.876
@@
DAV2-vits(Yang et al. 2024b) Least Squares 0.330 2.053 6.737 1.729 0.292 0.423 0.790 0.877
@@
DAV1-vits(Yang et al. 2024a) Least Squares 0.331 2.078 6.772 1.714 0.291 0.423 0.786 0.875
@@
Table 6: More detailed quantitative depth comparison on the KITTI dataset. Best results are in bold, second best are underlined.
@@
Table 7: Quantitative results on the NYUv2 dataset comparing VGLD and RSA in response to different textual descriptions. The LM_shift and LM_scale represent scalars values fitted using the Levenberg-Marquardt method. Best results are in bold.
@@
1 Text-1 RSA(Zeng et al. 2024b) 0.097 4.060 0.926 1.005 1.011
VGLD 0.075 3.562 0.949 1.004 1.001
Text-2 RSA(Zeng et al. 2024b) 0.084 4.251 0.923 1.007 1.010
VGLD 0.077 3.067 0.949 1.004 1.010
Text-3 RSA(Zeng et al. 2024b) 0.072 3.140 0.952 1.004 1.010
VGLD 0.068 3.088 0.951 1.004 1.003 1.010 1.010
2 Text-1 RSA(Zeng et al. 2024b) 0.108 2.327 0.905 1.009 1.014
VGLD 0.063 1.887 0.984 1.135 1.015
Text-2 RSA(Zeng et al. 2024b) 0.281 5.341 0.537 1.006 1.013
VGLD 0.099 2.144 0.915 1.010 1.014
Text-3 RSA(Zeng et al. 2024b) 0.109 2.157 0.906 1.011 1.014
VGLD 0.073 1.861 0.952 1.011 1.009 1.015 1.017
3 Text-1 RSA(Zeng et al. 2024b) 0.268 6.526 0.740 1.004 1.009
VGLD 0.119 2.516 0.919 1.009 1.010
Text-2 RSA(Zeng et al. 2024b) 0.171 4.479 0.854 1.005 1.010
VGLD 0.077 2.287 0.942 1.008 1.010
Text-3 RSA(Zeng et al. 2024b) 0.081 2.421 0.938 1.007 1.010
VGLD 0.062 2.236 0.953 1.009 1.008 1.011 1.011
Table 9: Quantitative results on the KITTI dataset comparing VGLD and RSA in response to different textual descriptions. The LM_shift and LM_scale represent scalars values fitted using the Levenberg-Marquardt method. Best results are in bold.
@@
RDE Model Method Lower is better Higher is better
ZoeDepth(Bhat et al. 2023) robust depth $\dagger$ 0.123 0.356 0.053 0.856 0.979 0.995
ScaleDepth(Zhu et al. 2024) 0.129 0.359 0.866
MiDas-1(Reiner et al. 2023) Least Squares 0.197 0.418 0.346 0.278 0.061 0.873 0.964 0.981
Levenberg Marquardt 0.158 0.440 0.252 0.116 0.032 0.950 0.988 0.995
RSA-NK(Zeng et al. 2024b) 0.299 0.589 0.575 0.251 0.094 0.615 0.900 0.977
VGLD-NK-T(Ours) 0.318 0.647 0.566 0.242 0.089 0.643 0.914 0.980
VGLD-NK-I(Ours) 0.259 0.595 0.468 0.202 0.071 0.751 0.957 0.991
VGLD-NK-TCI(Ours) 0.262 0.628 0.467 0.202 0.071 0.751 0.959 0.991
MiDas-2(Ranftl et al. 2020) Least Squares 0.203 0.419 0.365 0.272 0.062 0.860 0.961 0.981
Levenberg Marquardt 0.173 0.438 0.291 0.132 0.039 0.926 0.984 0.994
VGLD-NK-T(Ours) 0.316 0.795 0.597 0.249 0.090 0.639 0.908 0.978
VGLD-NK-I(Ours) 0.288 0.688 0.552 0.246 0.090 0.627 0.922 0.984
VGLD-NK-TCI(Ours) 0.275 0.670 0.513 0.225 0.080 0.694 0.941 0.987
DAV2-vits(Yang et al. 2024b) Least Squares 0.194 0.418 0.337 0.305 0.062 0.880 0.963 0.980
Levenberg Marquardt 0.146 0.439 0.224 0.103 0.027 0.961 0.989 0.995
VGLD-NK-T(Ours) 0.304 0.742 0.564 0.236 0.089 0.644 0.920 0.983
VGLD-NK-I(Ours) 0.273 0.564 0.535 0.236 0.090 0.617 0.931 0.989
VGLD-NK-TCI(Ours) 0.241 0.545 0.433 0.189 0.067 0.779 0.967 0.993
DAV1-vits(Yang et al. 2024a) Least Squares 0.196 0.416 0.341 0.282 0.061 0.875 0.963 0.981
Levenberg Marquardt 0.151 0.440 0.234 0.108 0.029 0.957 0.989 0.995
RSA-NK(Zeng et al. 2024b) 0.290 0.563 0.571 0.250 0.092 0.640 0.899 0.969
VGLD-NK-T(Ours) 0.281 0.583 0.532 0.214 0.078 0.711 0.945 0.987
VGLD-NK-I(Ours) 0.250 0.573 0.443 0.194 0.070 0.764 0.965 0.991
VGLD-NK-TCI(Ours) 0.241 0.545 0.433 0.189 0.067 0.779 0.967 0.993
Table 11: Zero-shot generalization to SUN-RGBD (Indoor). Best results are in bold, second best are underlined.